# A neutral-atom Hubbard quantum simulator in the cryogenic regime

### v1

@@ Abstract
Ultracold fermionic atoms in optical lattices offer pristine realizations of Hubbard models [1], which are fundamental to modern condensed matter physics and the study of strongly-correlated quantum materials [2, 3]. Despite significant advancements [4, 5, 6], the accessible temperatures in these optical lattice material analogs are still too high to address many open problems beyond the reach of current numerical techniques [7, 8, 9, 10]. Here, we demonstrate a several-fold reduction in temperature [11, 6, 12, 13], bringing large-scale quantum simulations of the Hubbard model into an entirely new regime. This is accomplished by transforming a low entropy product state into strongly-correlated states of interest via dynamic control of the model parameters [14, 15], which is extremely challenging to simulate classically and so explored using the quantum simulator itself [10]. At half filling, the long-range antiferromagnetic order is close to saturated, leading to a temperature of $T/t=0.05_{-0.05}^{0.06}$ based on comparisons to numerically exact simulations. Doped away from half-filling no unbiased numerical simulation is available [9]. Importantly, we are able to use quantum simulation to identify a new pathway for achieving similarly low temperatures with doping. This is confirmed by comparing short-range spin correlations to state-of-the-art, but approximate, constrained-path auxiliary field quantum Monte Carlo simulations [16, 17, 18]. Compared to the cuprates [19, 20, 2], the reported temperatures correspond to a reduction from far above to significantly below room temperature, where physics such as the pseudogap and stripe phases may be expected [19, 21, 22, 23, 3, 24]. Our work opens the door to quantum simulations that solve open questions in material science, develop synergies with numerical methods and theoretical studies, and lead to discoveries of new physics [8, 10].
@@ Introduction
The Hubbard model is a paradigmatic description of strongly-correlated electrons that is central to our understanding of quantum materials [3]. After decades of concentrated study, it is strongly believed that this model hosts a variety of exotic quantum states [25, 26], including the intriguing phases observed in the cuprate high-temperature superconductors [2, 22, 3]. Despite its apparent simplicity, solving the two-dimensional, square-lattice Hubbard model has proved an outstanding challenge. In recent years, building on decades of algorithmic development, combined use of the most advanced numerical methods has established certain aspects of the model’s properties [27, 21, 23, 24, 28]. Nevertheless, a full description of the low-temperature phase diagram remains beyond the reach of theoretical tools and computational methods [9].
@@
tp
tp
The BI can be prepared with extremely high fidelity because the chemical potential lies within the band gap featuring low density of states, which allows for efficient entropy redistribution to a metallic reservoir [41, 14, 11, 29, 34]. We load a spin-balanced mixture of fermionic 6Li atoms in the lowest two hyperfine states into an optical lattice (long-spacing lattice). We set the magnetic bias field at $550$G close to a broad Feshbach resonance to minimize interaction strengths for BI formation. Programmable optical potentials created by two separate Digital Micromirror Devices (DMDs) first confine about $340$ atoms into a BI covering $170$ sites, then isolate the BI from the reservoir. We estimate a low initial entropy per particle of $s=0.025(4)k_{B}$ [29] based on a measured singly-occupied site density of $n_{s}=0.7(2)\%$ in a central disk of $\simeq 110$ sites, although an accurate estimation is complicated by gradients at the edge of the confining potential due to finite optical resolution (see Methods).
Keeping the subsequent transformation slow relative to the relevant many-body timescales is crucial to realizing low final temperatures. This is particularly challenging as this protocol involves a substantial change in lattice filling from two in the BI to around one atom per site in strongly-correlated Hubbard systems. Such a massive change in the state typically involves very slow many-body timescales, which requires excellent local control of density [36, 29, 34, 37, 38]. To address this challenge, we take advantage of an optical lattice whose geometry is dynamically adjustable [43, 44, 12]. In particular, we continuously split single lattice sites into double-wells, and subsequently connect these double-wells together into a square lattice, realizing a scheme similar to the one proposed in [15]. This doubles the number of sites at fixed total atom number (Fig. Acknowledgementsc,d and Fig. .17), and naturally converts the BI into half-filled Fermi systems. However, this doubling is not exact in the experiment due to variations of the optical potentials. To precisely reach the target filling, our protocol further relies on tunable interactions and DMD potentials to ensure controlled preparation, isolation and expansion of the system of interest (see Methods).
@@
In the $\alpha\rightarrow 0$ limit of disconnected dimers, which corresponds to the BI, we only detect saturated spin correlations between nearest-neighbor sites, $|\mathbf{d}|=1$ within the dimers and consistent with 00 everywhere else (Fig. Acknowledgementsb). As these dimers are connected, spin correlations start to grow between the dimers at $\alpha\simeq 0.7$ and become uniform and isotropic in the square lattice $\alpha=1.0$.
@@
We experimentally check ramp adiabaticity and confirm that the spin correlations have saturated by varying the ramp duration, which could be limited by heating. Surprisingly, we find the strengths of the DMD potentials confining the Mott insulator, which may lead to atom transport into and out of the central region, can strongly affect the final spin correlations, and therefore temperatures (see Methods and Fig. .17). These measurements suggest charge transport may be crucial to temperature optimization [47, 48] in addition to spin dynamics of the idealized model.
@@
With the optimized experimental sequence, we observe antiferromagnetic spin correlations $C^{zz}_{\mathbf{d}}$ extending across the entire half-filling region with a radius $r=5$ for an interaction strength of $U/t=8.3(2)$ (Fig. Acknowledgementsc). We compare these experimentally measured correlations after sign-correction and azimuthal average $(-)^{|d_{x}+d_{y}|}C^{zz}_{d}$ to numerical results from simulations using finite-temperature determinant quantum Monte Carlo (DQMC) and ground-state auxiliary-field quantum Monte Carlo (AFQMC) methods, under open boundary conditions. At half-filling, the numerical results are in principle exact, although care must be taken to obtain accurate and unbiased spin correlations, especially at larger bond distances $d$ (see Methods). We find the measured $(-)^{|d_{x}+d_{y}|}C^{zz}_{d}$ are close to saturation and consistent with numerical data at $T/t\in[0,0.1]$ (Fig. Acknowledgementsd). These strong, long-ranged correlations translate to a narrow peak in the spin structure factor $S(\mathbf{q})$ at quasimomentum $\mathbf{q}=(\pi,\pi)$ (Fig. Acknowledgementse), which is obtained as the Fourier transformation of $C^{zz}_{\mathbf{d}}$ (see Methods). Comparing numerical DQMC data to the experimental value of the squared staggered magnetization $(m^{z})^{2}$ computed up to a cutoff $d_{\text{max}}=6$ allows us to extract a temperature of $T/t=0.05_{-0.05}^{0.06}$ (Fig. Acknowledgementsf and Fig. .17).
@@
The most intriguing and poorly understood physics of the Hubbard model resides in the doped regime. At low temperatures of $T/t<0.2$, and with intermediate interactions of $U/t\simeq 8$, the doped Hubbard model is extremely challenging. Unlike for half-filling, no computational method is available which can systematically provide exact numerical results for sufficiently large system sizes in this regime. Extending our cooling scheme to the above regime is therefore very valuable to advance our understanding, since quantum simulation can offer direct measurements, potentially adding a powerful new element in conjunction with approximate numerical simulations and analytic theory.
@@
Our cooling scheme for doped systems is thus as follows (Fig. Acknowledgementsa,b): after forming the BI, we reduce the lattice depth by $80\%$ and perform splitting. The large tunnelings and weak interactions transform the BI into a Fermi liquid. We then expand the atoms by reducing the strength of the confining DMD potential, while ramping the magnetic bias field to its final value of $590$G. The depth of the square lattice is then ramped up to its final value, reaching the strongly-correlated regime of $U/t=8.0(3)$. At the same time, the strengths of the DMD potentials are adjusted to minimize transport during lattice loading [47, 48] (see Methods).
@@
Interestingly, for bond distances of $d>1$, although the agreement is good on the scale of the variation between modestly low temperatures ($T/t=0.25$) and the ground state, the experiment shows stronger correlations than predicted by CP-AFQMC. It is not quite at the level of the agreement found at half-filling between experimental data and numerically exact results for all bond distances. We verify that our observations do not depend on the weak spatial variation of doping due to harmonic confinement by changing the radius $r$ of the experimental ROI, and further confirm that the system is at thermal equilibrium (see Fig. .17). Uncertainty in the calibration of experimental parameters (see Methods) is insufficient to explain the deviation, but other details in Hamiltonian difference (e.g. higher order corrections) could potentially contribute, especially at lower temperatures. The discrepancy may also hint at an underestimation of long-range correlations by CP-AFQMC. This would be consistent with data at elevated temperatures of $T/t=0.25,0.33$, where numerically exact DQMC results seem to be more consistent with experimentally measured spin correlations, and CP-AFQMC results show slightly lower correlations (see Methods and Fig. .17). These systematic comparisons are an example of the synergies that now exist between quantum simulations and numerics in the regime where exact numerical results are not available.
@@
In this work, we have demonstrated a substantial reduction in the temperatures achievable in large-scale quantum simulations of the Hubbard model, both with and without doping. Through careful dynamical tuning of parameters in a programmable optical lattice, we are able to transform a trivial low entropy state into a cold, strongly-correlated quantum many-body state in equilibrium. Critically, the dynamics involved in this transformation are infeasible to tackle via analytical techniques, or to simulate classically, and so must be optimized empirically using the quantum simulator itself. The above approach is broadly applicable to situations where one can prepare low entropy product states with high fidelity, for example with optical tweezers [51]. These techniques can also be extended to other lattice geometries, including triangular and kagome lattices, which may host quantum spin liquids [26], and square lattices with diagonal tunneling $t^{\prime}$, which may host unconventional superconductivity [52].
@@ Acknowledgements
f3 [Uncaptioned image] Cold Hubbard antiferromagnets in the strongly-correlated regime. (a), Intermediate interaction strengths, here $U/t=8.3(2)$, lead to density fluctuations in addition to spin fluctuations at half-filling. (b), To achieve the target filling (here $n=1$), we adjust the confining potential created by DMDs to spill excess atoms. A repulsive wall isolates the cold region from the reservoir, while a volcano-shaped potential enables a spilling of excess atoms from the center. (c), Spin correlations $C^{zz}_{\mathbf{d}}$ as a function of bond displacement $\mathbf{d}$, and averaged over a radius $r=5$ region, show nearly saturated long-range antiferromagnetic correlations. (d), Experimental data is compatible with numerically exact simulations (lines, DQMC and AFQMC) at temperatures $T/t<0.1$ and interactions $U/t=8$. (e), Low temperatures yield a sharp peak of the spin structure factor $S(\textbf{q})$ at quasi-momentum $\textbf{q}=(\pi,\pi)$. (f), We estimate temperature by computing the staggered magnetization with cutoff $d=6$ on bond distance, for both DQMC–AFQMC data (interpolated to $U/t=8.3(2)$, line) and experimental data ($1\sigma$ confidence interval shown as shaded area). Their comparison yields a temperature $T/t=0.05_{-0.05}^{0.06}$ (Methods).
@@
f4 [Uncaptioned image] Hole-doping cold Fermi Hubbard systems. (a),(b), To coherently introduce hole dopants into the half-filled state after splitting the BI, we increase the kinetic energy and decrease the interaction strength by reducing the lattice depths out of the tight-binding regime. This prepares cold weakly interacting Fermi liquid that inherits the low initial entropy, which is adiabatically expanded to reach a given target density. Reloading the lattices and ramping up interactions create cold hole-doped Hubbard systems at $U/t=8.0(3)$ (III). Schemes of splitting the BI into a half-filled antiferromagnetic Mott insulator at strong interaction of $U/t=18.6(8)$ (I) and intermediate interaction of $U/t=8.3(2)$ (II) are also marked. (c), Doping scheme in quasimomentum space. Splitting the BI, which completely fills the Brillouin zone, into the square lattice doubles the number of sites and the size of the Brillouin zone. The population of quasimomentum states remain nearly unchanged. Expanding in real space then decreases the size of the Fermi surface and coherently introduces doping. (d) Spin correlations as a function of bond displacements $C^{zz}_{\mathbf{d}}$ measured in an ROI of $r=3$. The range and magnitude of the antiferromagnetic correlation decreases with increased doping. (e) Azimuthal average of the sign-corrected spin correlations shown in (d). At small dopings of $\delta\leq 10\%$, the antiferromagnetism still remains long-ranged over the ROI. As we increase the number of hole dopants, the strengths of spin correlations at all distances are reduced by similar amounts. (f) Short-range spin correlations $C^{zz}_{d}$ at different distances $d=1,\sqrt{2},2,\sqrt{5}$ as we dope the system. The nearest-neighbor correlations at $d=1$ show quantitative agreement between experimental data and CP-AFQMC simulations at $U/t=8$ (solid line). At longer bond distances $d>1$, experimental data show stronger correlations than CP-AFQMC simulations.

### v2

# dfigure[$\NewEnvirondfigure*[1]$

@@ Abstract
Ultracold fermionic atoms in optical lattices offer pristine realizations of Hubbard models [1], which are fundamental to modern condensed matter physics [2, 3]. Despite significant advancements [4, 5, 6], the accessible temperatures in these optical lattice material analogs are still too high to address many open problems [7, 8, 9, 10]. Here, we demonstrate a several-fold reduction in temperature [11, 6, 12, 13], bringing large-scale quantum simulations of the Hubbard model into an entirely new regime. This is accomplished by transforming a low entropy product state into strongly-correlated states of interest via dynamic control of the model parameters [14, 15], which is extremely challenging to simulate classically [10]. At half filling, the long-range antiferromagnetic order is close to saturated, leading to a temperature of $T/t=0.05_{-0.05}^{0.06}$ based on comparisons to numerically exact simulations. Doped away from half-filling, it is exceedingly challenging to realize systematically accurate and predictive numerical simulations [9]. Importantly, we are able to use quantum simulation to identify a new pathway for achieving similarly low temperatures with doping. This is confirmed by comparing short-range spin correlations to state-of-the-art, but approximate, constrained-path auxiliary field quantum Monte Carlo simulations [16, 17, 18]. Compared to the cuprates [19, 20, 2], the reported temperatures correspond to a reduction from far above to below room temperature, where physics such as the pseudogap and stripe phases may be expected [19, 21, 22, 23, 3, 24]. Our work opens the door to quantum simulations that solve open questions in material science, develop synergies with numerical methods and theoretical studies, and lead to discoveries of new physics [8, 10].
@@ Introduction
The Hubbard model is a paradigmatic description of strongly-correlated electrons that is central to our understanding of quantum materials [3]. After decades of concentrated study, it is strongly believed that this model hosts a variety of exotic quantum states [25, 26], including the intriguing phases observed in the cuprate high-temperature superconductors [2, 22, 3]. Despite its apparent simplicity, solving the two-dimensional, square-lattice Hubbard model has proved an outstanding challenge. In recent years, building on decades of algorithmic development, combined use of the most advanced numerical methods has established certain aspects of the model’s properties [27, 21, 28, 29, 23, 30, 24]. Nevertheless, a full description of the low-temperature phase diagram remains beyond the reach of theoretical tools and computational methods [9].
@@
The BI can be prepared with extremely high fidelity because the chemical potential lies within the band gap featuring low density of states, which allows for efficient entropy redistribution to a metallic reservoir [43, 14, 11, 31, 36]. We load a spin-balanced mixture of fermionic 6Li atoms in the lowest two hyperfine states into an optical lattice (long-spacing lattice). We set the magnetic bias field at 550550G close to a broad Feshbach resonance to minimize interaction strengths for BI formation. Programmable optical potentials created by two separate Digital Micromirror Devices (DMDs) first confine about $340$ atoms into a BI covering $170$ sites, then isolate the BI from the reservoir. We estimate a low initial entropy per particle of $s=0.025(4)k_{B}$ [31] based on a measured singly-occupied site density of $n_{s}=0.7(2)\%$ in a central disk of $\simeq 110$ sites, although an accurate estimation is complicated by gradients at the edge of the confining potential due to finite optical resolution (see Methods).
Keeping the subsequent transformation slow relative to the relevant many-body timescales is crucial to realizing low final temperatures. This is particularly challenging as this protocol involves a substantial change in lattice filling from two in the BI to around one atom per site in strongly-correlated Hubbard systems. Such a massive change in the state typically involves very slow many-body timescales, which requires excellent local control of density [38, 31, 36, 39, 40]. To address this challenge, we take advantage of an optical lattice whose geometry is dynamically adjustable [44, 45, 12]. In particular, we continuously split single lattice sites into double-wells, and subsequently connect these double-wells together into a square lattice, realizing a scheme similar to the one proposed in [15]. This doubles the number of sites at fixed total atom number (Fig. Acknowledgementsc,d and Extended Data Fig. Competing Interests), and naturally converts the BI into half-filled Fermi systems. However, this doubling is not exact in the experiment due to variations of the optical potentials. To precisely reach the target filling, our protocol further relies on tunable interactions and DMD potentials to ensure controlled preparation, isolation and expansion of the system of interest (see Methods).
@@
In the $\alpha\rightarrow 0$ limit of disconnected dimers, which corresponds to the BI, we only detect saturated spin correlations between nearest-neighbor sites, $|\mathbf{d}|=1$ within the dimers and consistent with 0 everywhere else (Fig. Acknowledgementsb). As these dimers are connected, spin correlations start to grow between the dimers at $\alpha\simeq 0.7$ and become uniform and isotropic in the square lattice $\alpha=1.0$.
@@
We experimentally check ramp adiabaticity and confirm that the spin correlations have saturated by varying the ramp duration, which could be limited by heating. Surprisingly, we find the strengths of the DMD potentials confining the Mott insulator, which may lead to atom transport into and out of the central region, can strongly affect the final spin correlations, and therefore temperatures (see Methods and Extended Data Fig. Competing Interests). These measurements suggest charge transport may be crucial to temperature optimization [49, 50] in addition to spin dynamics of the idealized model.
@@
With the optimized experimental sequence, we observe antiferromagnetic spin correlations $C^{zz}_{\mathbf{d}}$ extending across the entire half-filling region with a radius $r=5$ for an interaction strength of $U/t=8.3(2)$ (Fig. Acknowledgementsc, see Methods). We compare these experimentally measured correlations after sign-correction and azimuthal average $(-)^{|d_{x}+d_{y}|}C^{zz}_{d}$ to numerical results from simulations using finite-temperature determinant quantum Monte Carlo (DQMC) and ground-state auxiliary-field quantum Monte Carlo (AFQMC) methods, under open boundary conditions. At half-filling, the numerical results are in principle exact, although care must be taken to obtain accurate and unbiased spin correlations, especially at larger bond distances $d$ (see Methods). We find the measured $(-)^{|d_{x}+d_{y}|}C^{zz}_{d}$ are close to saturation and consistent with numerical data at $T/t\in[0,0.1]$ (Fig. Acknowledgementsd). These strong, long-ranged correlations translate to a narrow peak in the spin structure factor $S(\mathbf{q})$ at quasimomentum $\mathbf{q}=(\pi,\pi)$ (Fig. Acknowledgementse), which is obtained as the Fourier transformation of $C^{zz}_{\mathbf{d}}$ (see Methods). Comparing numerical DQMC data to the experimental value of the squared staggered magnetization $(m^{z})^{2}$ computed up to a cutoff $d_{\text{max}}=6$ allows us to extract a temperature of $T/t=0.05_{-0.05}^{0.06}$ (Fig. Acknowledgementsf and Extended Data Fig. Competing Interests).
@@
The most intriguing and poorly understood physics of the Hubbard model resides in the doped regime. At low temperatures of $T/t<0.2$, and with intermediate interactions of $U/t\simeq 8$, the doped Hubbard model is extremely challenging. Unlike for half-filling, no numerically exact results for spin correlations $C^{zz}_{d}$ are available in this regime for sufficiently large system sizes, which have systematically explored parameters such as temperature and interaction. Extending our cooling scheme to the above regime is therefore very valuable to advance our understanding, since quantum simulation can offer direct measurements, potentially adding a powerful new element in conjunction with approximate numerical simulations and analytic theory.
@@
Our cooling scheme for doped systems is thus as follows (Fig. Acknowledgementsa,b): after forming the BI, we reduce the lattice depth by $80\%$ and perform splitting. The large tunnelings and weak interactions transform the BI into a Fermi liquid. We then expand the atoms by reducing the strength of the confining DMD potential, while ramping the magnetic bias field to its final value of 590590G. The depth of the square lattice is then ramped up to its final value, reaching the strongly-correlated regime of $U/t=8.0(3)$ (see Methods). At the same time, the strengths of the DMD potentials are adjusted to minimize transport during lattice loading [49, 50] (see Methods).
@@
Interestingly, for bond distances of $d>1$, although the agreement is good on the scale of the variation between modestly low temperatures ($T/t=0.25$) and the ground state, the experiment shows stronger correlations than predicted by CP-AFQMC. It is not quite at the level of the agreement found at half-filling between experimental data and numerically exact results for all bond distances. We verify that our observations do not depend on the weak spatial variation of doping due to harmonic confinement by changing the radius $r$ of the experimental ROI, and further confirm that the system is at thermal equilibrium (see Extended Data Fig. Competing Interests). Uncertainty in the calibration of experimental parameters (see Methods) is insufficient to explain the deviation, but other details in Hamiltonian difference (e.g. higher order corrections) could potentially contribute, especially at lower temperatures. The discrepancy may also hint at an underestimation of long-range correlations by CP-AFQMC. This would be consistent with data at elevated temperatures of $T/t=0.25,0.33$, where numerically exact DQMC results seem to be more consistent with experimentally measured spin correlations, and CP-AFQMC results show slightly lower correlations (see Methods and Extended Data Fig. Competing Interests). These systematic comparisons are an example of the synergies that now exist between quantum simulations and numerics in the regime where obtaining exact numerical results are incredibly challenging.
@@
In this work, we have demonstrated a substantial reduction in the temperatures achievable in large-scale quantum simulations of the Hubbard model, both with and without doping. Through careful dynamical tuning of parameters in a programmable optical lattice, we are able to transform a trivial low entropy state into a cold, strongly-correlated quantum many-body state in equilibrium. Critically, the dynamics involved in this transformation are exceedingly difficult to tackle via analytical techniques, or to simulate classically, and so must be optimized empirically using the quantum simulator itself. The above approach is broadly applicable to situations where one can prepare low entropy product states with high fidelity, for example with optical tweezers [53]. These techniques can also be extended to other lattice geometries, including triangular and kagome lattices, which may host quantum spin liquids [26], and square lattices with diagonal tunneling $t^{\prime}$, which may host unconventional superconductivity [54].
@@ Acknowledgements
f3 [Uncaptioned image] Cold Hubbard antiferromagnets in the strongly-correlated regime. (a), Intermediate interaction strengths, here $U/t=8.3(2)$, lead to density fluctuations in addition to spin fluctuations at half-filling. (b), To achieve the target filling (here $n=1$), we adjust the confining potential created by DMDs to spill excess atoms. A repulsive wall isolates the cold region from the reservoir, while a volcano-shaped potential enables a spilling of excess atoms from the center. (c), Spin correlations $C^{zz}_{\mathbf{d}}$ as a function of bond displacement $\mathbf{d}$, and averaged over a radius $r=5$ region, show nearly saturated long-range antiferromagnetic correlations. The error bars denote one s.e.m. (d), Experimental data is compatible with numerically exact simulations (lines, DQMC and AFQMC) at temperatures $T/t<0.1$ and interactions $U/t=8$. (e), Low temperatures yield a sharp peak of the spin structure factor $S(\textbf{q})$ at quasi-momentum $\textbf{q}=(\pi,\pi)$. (f), We estimate temperature by computing the staggered magnetization with cutoff $d=6$ on bond distance, for both DQMC–AFQMC data (interpolated to $U/t=8.3(2)$, line) and experimental data ($1\sigma$ confidence interval shown as shaded area). Their comparison yields a temperature $T/t=0.05_{-0.05}^{0.06}$ (Methods).
@@
f4 [Uncaptioned image] Hole-doping cold Fermi Hubbard systems. (a),(b), To coherently introduce hole dopants into the half-filled state after splitting the BI, we increase the kinetic energy and decrease the interaction strength by reducing the lattice depths out of the tight-binding regime. This prepares cold weakly interacting Fermi liquid that inherits the low initial entropy, which is adiabatically expanded to reach a given target density. Reloading the lattices and ramping up interactions create cold hole-doped Hubbard systems at $U/t=8.0(3)$ (III). Schemes of splitting the BI into a half-filled antiferromagnetic Mott insulator at strong interaction of $U/t=18.6(8)$ (I) and intermediate interaction of $U/t=8.3(2)$ (II) are also marked. (c), Doping scheme in quasimomentum space. Splitting the BI, which completely fills the Brillouin zone, into the square lattice doubles the number of sites and the size of the Brillouin zone. The population of quasimomentum states remain nearly unchanged. Expanding in real space then decreases the size of the Fermi surface and coherently introduces doping. (d) Spin correlations as a function of bond displacements $C^{zz}_{\mathbf{d}}$ measured in an ROI of $r=3$. The range and magnitude of the antiferromagnetic correlation decreases with increased doping. (e) Azimuthal average of the sign-corrected spin correlations shown in (d). At small dopings of $\delta\leq 10\%$, the antiferromagnetism still remains long-ranged over the ROI. As we increase the number of hole dopants, the strengths of spin correlations at all distances are reduced by similar amounts. (f) Short-range spin correlations $C^{zz}_{d}$ at different distances $d=1,\sqrt{2},2,\sqrt{5}$ as we dope the system. The nearest-neighbor correlations at $d=1$ show quantitative agreement between experimental data and CP-AFQMC simulations at $U/t=8$ and $T/t\in[0,0.1]$ (solid line). At longer bond distances $d>1$, experimental data show stronger correlations than CP-AFQMC simulations. The error bars denote one s.e.m.